with the tailwind a bird flew at ground speed of 3 mph flying the same path against the same wind the bird travels at a ground speed of 1.5 mph what is the birds airspeed what is the wind speed
I have s=3+t but I just don't understand how to write the equations for systems of equations

Answer by ankor@dixie-net.com(22740)

with the tailwind a bird flew at ground speed of 3 mph flying the same path against the same wind the bird travels at a ground speed of 1.5 mph
what is the birds airspeed
what is the wind speed
:
Let w = the speed of the wind
Let s = the speed of the bird in still air
then
(s+w) = effective speed with the wind
and
(s-w) = effective speed against the wind
:
Write an equation for each way
s + w = 3
s - w = 1.5
--------------addition eliminates w find s
2s = 4.5
s = 4.5/2
s = 2.25 mph is the bird airspeed
:
Find the wind using the 1st equation
2.25 + w = 3
w = 3 - 2.25
w = .75 mph is the speed of the wind
:
:
Check these solutions in the 2nd equation
2.25 - .75 = 1.5